OpenStudy (anonymous):

(6z^2-4z+1)(8-3z) @uri 48z^2-18z+8?

OpenStudy (uri):

Can you show me the steps how did you do it? @TheGstudent23

OpenStudy (anonymous):

6z^2 x 8=48z^2 6z^2 x -3z=24z^3 -4z x 8=-32z -4z x -12z^2 1x8=8 1 x -3z=3z

OpenStudy (uri):

Continue.

OpenStudy (anonymous):

48z^2-12z^2=36z^2+24^3-32z+8 i have no idea what i did the first time

OpenStudy (uri):

6 x-3=18 not 24. Anyway, 6z^2(8z-3z) -4z(8-3z) +1(8-3z) 48x^2-18z^3-32z+12z^2+8-3z

OpenStudy (uri):

Now solve this,bring all values of samew exponent together.

OpenStudy (anonymous):

-18z^3+60z^2-35z+35z+8

OpenStudy (anonymous):

sorry forgot a lot about multiplication

OpenStudy (anonymous):

-18z^3+60z^2-35z+8 typo

OpenStudy (uri):

@TheGstudent23 Always go step by step you missed this step.. 48x^2-18z^3-32z+12z^2+8-3z 48x^2+12x^2-18x^3-32z-3z+8 Then this, -18z^3+60z^2-35z+8

OpenStudy (anonymous):

ok thank you

OpenStudy (uri):

You can lose marks if you don't go step by step even if your answer is right..